# Bayesian Prediction under Moment Conditioning

### v1

@@ 3.3 Tilted Finite Gaussian de Finetti Theorem
Let $\phi_{H^{*}}(v)=|H^{*}|^{1/2}(2\pi)^{-r/2}\exp\!\big(-\tfrac{1}{2}v^{\top}H^{*}v\big)$ denote the Gaussian density on $\mathsf{T}^{*}$. For measurable $A\subset\mathcal{X}^{m}$ with $m=O(n^{\eta})$ for any fixed $\eta<1/2$,
@@
where $P(v)=\Pi(P^{*}+n^{-1/2}v)$ is the local chart of $E$, and
@@
Conditioning on the constraints replaces the flat multinomial prior with a Gaussian measure on the manifold, centred at $P^{*}$ with covariance $(nH^{*})^{-1}$. Predictive collapse is thus a geometric manifestation of curvature: steep curvature compresses the Gaussian, whereas flat curvature leaves it diffuse. This provides a finite, curvature-adaptive analogue of de Finetti’s representation for exchangeable sequences.
@@
The mixture in Theorem 3.1 is not uniform over empirical types but exponentially tilted around the information projection $P^{*}$. The quadratic term $v^{\top}H^{*}v$ encodes the curvature of the constraint manifold, so the resulting representation is a tilted finite de Finetti theorem: exchangeability holds, but the prior over empirical laws is weighted by constraint-induced information curvature. This tilt unifies the de Finetti mixture with exponential weighting, providing the finite predictive counterpart to large-deviation geometry.
@@ 3.4 Curvature and Predictive Contraction
The Gaussian approximation implies that deviations of the constrained predictive law $\mu_{n,m}$ from its deterministic limit $(P^{*})^{\otimes m}$ arise from three finite-sample sources. The first is the hypergeometric correction induced by sampling without replacement, which contributes a term of order $O(m^{2}/(np_{\min}^{*}))$. The second is local deviation within the Lanford window, controlled by the Lipschitz continuity of the likelihood and scaling as $O(m\rho_{n}n^{-1/2})$. The third is the exponentially small tail mass outside the window, of order $\exp(-\tfrac{1}{2}\lambda_{\min}\rho_{n}^{2})=n^{-1}$. Balancing the first two effects by the optimal choice $\rho_{n}^{2}=2\log n/\lambda_{\min}(H^{*})$ gives the final curvature-dependent rate.
Theorem 3.4 (Predictive collapse bound).
Under the conditions of Theorem 3.1
if $m=o\!\big(\sqrt{n\,\lambda_{\min}(H^{*})/\log n}\big)$, then
$\|\mu_{n,m}-(P^{*})^{\otimes m}\|_{\mathrm{TV}}\;\leq\;C_{\mathrm{geo}}\,m\sqrt{\frac{\log n}{n\,\lambda_{\min}(H^{*})}}\;+\;C^{\prime}_{\mathrm{geo}}\,\frac{m^{2}}{n\,p_{\min}^{*}}\;+\;o(n^{-1/2}),$
where $C_{\mathrm{geo}}$ and $C^{\prime}_{\mathrm{geo}}$ depend only on $(|\mathcal{X}|,d,p_{\min}^{*})$ and the local curvature of the constraint manifold.
@@
The first term represents geometric contraction: predictive uncertainty shrinks at rate $O\{(n\lambda_{\min}(H^{*}))^{-1/2}\}$, reflecting the curvature radius of the manifold. The second term captures finite-population correction, negligible whenever $m\ll\sqrt{n}$. Hence, stronger curvature or larger samples force the predictive law to concentrate sharply, its randomness collapsing onto $P^{*}$.
@@ 4 Asymptotic and Geometric Connections
The discrete Gaussian approximation developed above implies that constrained predictive laws satisfy a local asymptotic normality property analogous to that of parametric experiments. In the limit, the entire predictive mechanism behaves as a curved exponential family centered at the information projection $P^{*}$. This identification aligns with the differential-geometric structure of exponential families formalized by Brown [9] and connects to the asymptotic theory of statistical experiments initiated by Le Cam [35] and its modern expositions by van der Vaart [54], Le Cam and Yang [37], in which local asymptotic normality (LAN) provides the bridge between finite models and their Gaussian tangent experiments.
@@
Let $\Delta_{n}=\sqrt{n}\,(P_{n}-P^{*})$ and let $\xi_{m}$ denote the Gaussian shift induced by the tangent coordinates of the predictive experiment. Then, uniformly over bounded sequences of alternatives,
@@
Equation (4.1) identifies the constrained predictive mechanism with a Gaussian shift experiment on the tangent space $(\mathsf{T}^{*},H^{*})$. In Le Cam’s metric for statistical experiments, this means that the sequence $\{\mu_{n,m}\}$ converges to a limit experiment given by observation of a Gaussian random element with covariance $(H^{*})^{-1}$. The predictive collapse phenomenon, therefore, admits a coordinate-free formulation: it is precisely the convergence of predictive experiments to their Gaussian tangent model, with curvature determining the information content of the limit.
@@
The quantities appearing in our bounds are invariant under smooth reparameterizations of the constraint manifold. The matrix $H^{*}$ is the restriction of the Riemannian information metric to $\mathsf{T}^{*}$, and $\lambda_{\min}(H^{*})$ depends only on the intrinsic geometry of $E$, not on its coordinate representation. Thus, the results describe geometric, not coordinate-features of predictive concentration.
@@ 5 Connections with Exponentially Tilted Empirical Likelihood
Exponentially tilted empirical likelihood (ETEL) and its Bayesian counterpart (BETEL) have achieved notable success in semiparametric inference [48, 50, 27, 30], yet their theoretical justification is often presented as a pragmatic extension of empirical likelihood rather than as a direct consequence of fundamental probabilistic principles. Our framework derives ETEL and BETEL from first principles, showing that they arise naturally as conditional likelihoods under exchangeability and moment constraints. In particular, BETEL coincides with the conditional likelihood in the limit experiment defined by the constraint manifold.
@@
$\pi(\theta|X_{1:n})\propto\pi(\theta)\prod_{i=1}^{n}P_{\theta}^{*}(X_{i})\exp\!\Big(\lambda_{\theta}^{\top}\sum_{i=1}^{n}h(X_{i})\Big)(1+o_{p}(1)).$
Hence BETEL coincides with the conditional likelihood implied by exchangeability and moment constraints in the LAN limit experiment.
The proof follows from Theorem 3.1: conditioning on $P_{n}\in E_{\theta,n}$ confines empirical types to a Gaussian neighborhood of $P_{\theta}^{*}$, and Lagrange duality converts this weighting into the exponential tilt. This geometric interpretation explains several known properties. First, semiparametric efficiency arises because $P_{\theta}^{*}$ minimizes Kullback–Leibler divergence, so its Hessian $H_{\theta}^{*}$ provides the semiparametric information bound. Second, the Bernstein–von Mises behavior established by Chib and Simoni [11, 12] follows from the discrete Gaussian law: the BETEL posterior is asymptotically normal with covariance $H_{\theta_{0}}^{*-1}$. Third, over-identification robustness emerges naturally because additional valid moments enlarge curvature, increasing $\lambda_{\min}(H_{\theta}^{*})$ and hence accelerating contraction.
@@
Thus, BETEL inherits both its efficiency and its rates directly from the constraint manifold’s curvature, establishing it as the canonical Bayesian method for moment-restricted inference.
6 GMM and GEE as Quadratic Approximations
While BETEL provides the exact conditional likelihood, computational considerations often favor quadratic approximations. We show that generalized method of moments (GMM; 23) and generalized estimating equations (GEE; 39) arise naturally as optimal second-order expansions of this likelihood, with their properties governed by the same curvature parameter. This connection situates our framework within the broader theory of moment-based estimation and quasi-likelihood, linking the geometric weighting implied by $H^{*}$ to classical results on optimal GMM weighting and efficiency [41, 55, 26]. The curvature perspective thus provides a unified explanation for both efficiency under correct specification and robustness under model misspecification.
6.1 GMM as Optimal Quadratic Expansion
@@ 6.2 GEE Robustness via Effective Curvature
Hence, GEE achieves robustness by replacing the exact information metric with a working approximation. Independence working correlations preserve rates whenever true correlations are moderate, while pathological choices that nearly annihilate directions in $\mathrm{span}(D_{i})$ destroy stability by driving $\lambda_{\min}(J)$ to zero.
@@ 7.1 Berk’s Theorem via Constraint Curvature
This refines the misspecified Bernstein–von Mises picture [e.g. 31]: curvature dictates the local information content and the rate at which the posterior contracts around the pseudo-true parameter. Large curvature accelerates learning yet increases brittleness under model error; small curvature slows learning but yields thicker, more robust posteriors. The same phenomenon holds predictively.
@@
Thus, forecasts converge to the pseudo-true predictive law at the same curvature-controlled rate.
@@ 7.2 Safe Bayes and Geometric Tempering
Tempering, therefore, trades speed for robustness in a controlled, metric-level manner.
@@
Let $\widehat{\lambda}_{\min}$ be a consistent local estimate of $\lambda_{\min}(H^{*})$. Any choice $T=T(\widehat{\lambda}_{\min})$ that keeps $T/\widehat{\lambda}_{\min}$ bounded above and away from zero yields a collapse rate $O\!\big(m\sqrt{\log n/n}\big)$ up to constants uniform over models with comparable curvature. In particular, setting
@@
Implications. Berk’s theorem is not an exception to Bayesian learning but the boundary case of constrained prediction, with curvature serving as the single control parameter for both efficiency and robustness. The diagnostic is geometric (inspect $\lambda_{\min}(H^{*})$); the remedy is geometric (rescale the metric via tempering); and the consequences are transparent: fast and brittle when curvature is large, slow and robust when curvature is small, and optimally calibrated when curvature is moderated. This unifies misspecification analysis, posterior tempering, and predictive calibration within a single geometric framework.
@@ 8 Discussion and Outlook
Chernozhukov and Hong [10] reinterpret classical estimators such as GMM through quasi-posteriors, showing that standard estimators emerge as Laplace-type limits of MCMC-based Bayesian updates. Their framework uses MCMC to approximate optimal weighting and curvature. By contrast, the present analysis derives this Gaussian structure directly from first principles, exchangeability, and constraint geometry, without recourse to sampling algorithms. The curvature term $\lambda_{\min}(H^{*})$ appears not as a numerical by-product but as the intrinsic geometric rate governing predictive contraction, providing a constructive explanation of the phenomena that quasi-Bayesian methods simulate.
@@ 8.1 Methodological Implications
Our analysis complements the classical Sanov route by asking the dual question: given empirical constraints, how rapidly does the predictive law return to equilibrium on the constraint manifold? This perspective makes $\lambda_{\min}(H^{*})$ the control parameter for this return, linking information geometry with predictive concentration. The Lanford window construction refines standard saddlepoint expansions by balancing lattice discretization and Gaussian localization, yielding practical non-asymptotic bounds. As the construction is coordinate-invariant, the contraction rate depends only on geometric curvature, not on the parameterization.
@@ Acknowledgement
The initial idea for this work originated in 2014 after the authors attended a seminar on Bayesian nonparametrics and the method of moments. It was during that discussion that the possibility of an alternative route, linking predictive asymptotics to constraint geometry, first became apparent. The present framework was developed gradually from that early intuition.
@@ Appendix A Proofs of Main Results
Lanford Windows and Operator-Theoretic Foundation
The Lanford window construction in Section 3 arises from an operator-theoretic formulation originally developed by Lanford [33, 34] for lattice gases, here adapted to constrained exchangeable sequences.
The Gaussian approximation holds for $O(n^{-1/2})$ fluctuations, while the multinomial lattice contributes discretization errors of comparable order. Setting
@@
balances these effects: tails beyond $\rho_{n}$ carry mass $O(n^{-1})$, while lattice errors within $\mathcal{W}_{n}$ remain $O(n^{-1/2})$.
For empirical measures on $E_{n}$, define the Lanford operator
@@
which maps test functions to their Gaussian-weighted local averages around $P^{*}$.
The expected localization radius then satisfies
@@
This admits a unique fixed point with
@@
yielding
$\rho_{n}=\big(2\log n/\lambda_{\min}(H^{*})\big)^{1/2}.$
The window is therefore curvature-adaptive: it expands when the manifold is flat (small $\lambda_{\min}$) and contracts under strong curvature.
Iterating $\mathcal{L}_{n}$ contracts any initial radius estimate toward this fixed point at rate $1-O(n^{-1/2})$, ensuring a self-consistent localization that guarantees uniform Gaussian validity within $\mathcal{W}_{n}$.
A.1 Proof of Theorem 3.1
Decompose the predictive sum into interior and exterior contributions,
@@
where $S_{n}^{(\mathrm{in})}(A)$ aggregates terms with $P\in\mathcal{W}_{n}$ and $S_{n}^{(\mathrm{out})}(A)$ those with $P\notin\mathcal{W}_{n}$. The following lemma bounds the latter contribution.
@@
For $P\notin\mathcal{W}_{n}$,
@@
so that $|S_{n}^{(\mathrm{out})}(A)|=O(n^{-1})$ uniformly in $A\subset\mathcal{X}^{m}$.
Inside $\mathcal{W}_{n}$, the Heath–Sudderth mixture weights admit the local expansion
@@
where $P=P^{*}+n^{-1/2}v+O(n^{-1})$, $L_{n}(v;x_{1:m})=n^{-1/2}\sum_{i=1}^{m}v^{\top}\nabla\!\log P^{*}(x_{i})$, and $|R_{n}(v;m)|=O(\log n/n^{1/2})$ uniformly for $\|v\|\leq\rho_{n}$.
@@
Let $\mathcal{L}_{n}=\{\,n^{1/2}(P-P^{*}):P\in E_{n}\cap\mathcal{W}_{n}\,\}$. Then, for all functions $f$ with Gaussian envelopes,
@@
Evaluating the integral yields the Gaussian law on $\mathsf{T}^{*}$ with covariance $(H^{*})^{-1}$ and uniform approximation error $O(n^{-1/2}\operatorname{polylog}n)$. Combining Lemmas A.1 and A.2 therefore establishes Theorem 3.1.

### v2

@@ 3.3 Tilted Finite Gaussian de Finetti Theorem
Let
$\phi_{H^{*}}(v)=|H^{*}|^{1/2}(2\pi)^{-r/2}\exp\!\big(-\tfrac{1}{2}v^{\top}H^{*}v\big)$
denote the Gaussian density on the tangent space $\mathsf{T}^{*}$, where $r=\dim(\mathsf{T}^{*})$. For any measurable set $A\subset\mathcal{X}^{m}$ with $m=O(n^{\eta})$ for some fixed $\eta<1/2$,
@@
where $P(v)=\Pi(P^{*}+n^{-1/2}v)$ is the local chart projection of $P^{*}+n^{-1/2}v$ onto $E$, and the error term satisfies
@@
Conditioning on the constraints replaces the flat multinomial prior with a Gaussian measure on the manifold, centred at $P^{*}$ with covariance $(nH^{*})^{-1}$. Predictive collapse is therefore a geometric manifestation of curvature: steep curvature compresses the Gaussian, whereas flat curvature leaves it diffuse. The result provides a finite, curvature–adaptive analogue of de Finetti’s representation for exchangeable sequences.
@@
The mixture in Theorem 3.1 is not uniform over empirical types but exponentially tilted around the information projection $P^{*}$. The quadratic form $v^{\top}H^{*}v$ encodes the curvature of the constraint manifold, so the resulting representation is a tilted finite de Finetti theorem: exchangeability holds, but the prior over empirical laws is reweighted by the constraint–induced information curvature. This tilt unifies the de Finetti mixture with exponential weighting, providing the finite predictive counterpart to large–deviation geometry.
@@ 3.4 Curvature and Predictive Contraction
The Gaussian approximation of Theorem 3.1 implies that deviations of the constrained predictive law $\mu_{n,m}$ from its deterministic limit $(P^{*})^{\otimes m}$ arise from three finite–sample effects. The first is the hypergeometric correction due to sampling without replacement, contributing a term of order $O\!\{m^{2}/(np_{\min}^{*})\}$, where $p_{\min}^{*}=\min_{x}P^{*}(x)$. The second is the local deviation within the Lanford window $\mathcal{W}_{n}=\{P:\|P-P^{*}\|\leq\rho_{n}/\sqrt{n}\}$, controlled by the Lipschitz continuity of the likelihood and scaling as $O\!\bigl(m\,\rho_{n}/\sqrt{n}\bigr)$. The third is the exponentially small tail mass outside the window, of order $\exp\{-\tfrac{1}{2}\lambda_{\min}(H^{*})\rho_{n}^{2}\}=n^{-1}$. Balancing the first two effects yields the curvature–adaptive choice
$\rho_{n}^{2}=\frac{2\log n}{\lambda_{\min}(H^{*})}.$
Theorem 3.4 (Predictive collapse).
Assume the conditions of Theorem 3.1 and let $m=o\!\bigl(\sqrt{n\,\lambda_{\min}(H^{*})/\log n}\,\bigr)$. Then
$\bigl\|\mu_{n,m}-(P^{*})^{\otimes m}\bigr\|_{\mathrm{TV}}\;\leq\;C_{\mathrm{geo}}\,m\sqrt{\tfrac{\log n}{n\,\lambda_{\min}(H^{*})}}\;+\;C^{\prime}_{\mathrm{geo}}\,\tfrac{m^{2}}{n\,p_{\min}^{*}}\;+\;o(n^{-1/2}),$
where $C_{\mathrm{geo}}$ and $C^{\prime}_{\mathrm{geo}}$ depend only on $(|\mathcal{X}|,d,p_{\min}^{*})$ and on the local curvature of the constraint manifold near $P^{*}$.
@@
The leading term represents geometric contraction: predictive uncertainty shrinks at rate $O\bigl((n\,\lambda_{\min}(H^{*}))^{-1/2}\bigr)$, governed by the curvature radius of the constraint manifold. The second term is the finite–population correction, negligible whenever $m\ll\sqrt{n}$. Hence increasing curvature or enlarging the sample size forces the predictive law to collapse sharply onto the information projection $P^{*}$.
@@ 4 Asymptotic and Geometric Connections
The discrete Gaussian approximation developed above implies that constrained predictive laws satisfy a local asymptotic normality property analogous to that of parametric experiments. In the limit, the entire predictive mechanism behaves as a curved exponential family centred at the information projection $P^{*}$. This identification aligns with the differential-geometric structure of exponential families formalized by Brown [9] and connects to the asymptotic theory of statistical experiments initiated by Le Cam [39] and its modern expositions by van der Vaart [58], Le Cam and Yang [40], in which local asymptotic normality (LAN) provides the bridge between finite models and their Gaussian tangent experiments.
@@
Under the conditions of Theorem 3.1, let $\Delta_{n}=\sqrt{n}\,(P_{n}-P^{*})$ and let $\xi_{m}$ denote the Gaussian shift induced by the tangent coordinates of the predictive experiment. Then, uniformly over bounded sequences of alternatives,
@@
The LAN expansion in Corollary 4.1 identifies the constrained predictive mechanism with a Gaussian shift experiment on the tangent space $(\mathsf{T}^{*},H^{*})$. In Le Cam’s metric for statistical experiments, this means that the sequence $\{\mu_{n,m}\}$ converges to a limit experiment given by observation of a Gaussian random element with covariance $(H^{*})^{-1}$. The predictive collapse phenomenon therefore admits a coordinate-free formulation: it is precisely the convergence of predictive experiments to their Gaussian tangent model, with curvature determining the information content of the limit.
@@
The quantities appearing in our bounds are invariant under smooth reparameterizations of the constraint manifold. The matrix $H^{*}$ is the restriction of the Riemannian information metric to $\mathsf{T}^{*}$, and $\lambda_{\min}(H^{*})$ depends only on the intrinsic geometry of $E$, not on its coordinate representation. Thus the results describe geometric, not coordinate–features of predictive concentration.
@@ 5 Connections with Exponentially Tilted Empirical Likelihood
Exponentially tilted empirical likelihood (ETEL) and its Bayesian analogue (BETEL) have achieved notable success in semiparametric inference [52, 27, 32, 55, 38, 37]. Yet their theoretical justification is often presented as a pragmatic extension of empirical likelihood rather than as a direct consequence of probabilistic first principles. Our framework derives ETEL and BETEL from such principles, showing that they emerge naturally as conditional likelihoods under exchangeability and moment constraints. In particular, BETEL coincides with the conditional likelihood in the limit experiment defined by the curvature of the constraint manifold, thus linking de Finetti coherence with information-geometric inference.
@@
$\pi(\theta\mid X_{1:n})\propto\pi(\theta)\prod_{i=1}^{n}P_{\theta}^{*}(X_{i})\,(1+o_{p}(1)),$
where $P_{\theta}^{*}(x)=Q(x)\exp\{\lambda_{\theta}^{\top}h(x)\}/Z(\lambda_{\theta})$ already incorporates the exponential tilt. Hence BETEL coincides with the conditional likelihood implied by exchangeability and moment constraints in the LAN limit experiment.
The proof follows from Theorem 3.1: conditioning on $P_{n}\in E_{\theta,n}$ confines empirical types to a Gaussian neighborhood of $P_{\theta}^{*}$, and Lagrange duality converts this weighting into the exponential tilt. This geometric interpretation explains several known properties. First, semiparametric efficiency arises because $P_{\theta}^{*}$ minimizes Kullback–Leibler divergence, so its Hessian $H_{\theta}^{*}$ provides the semiparametric information bound. Second, the Bernstein–von Mises behavior established by Chib et al. [11, 12] follows from the discrete Gaussian law: the BETEL posterior is asymptotically normal with covariance $H_{\theta_{0}}^{*-1}$. Third, over-identification robustness emerges naturally because additional valid moments enlarge curvature, increasing $\lambda_{\min}(H_{\theta}^{*})$ and hence accelerating contraction.
@@
Thus BETEL inherits both its efficiency and its rates directly from the constraint manifold’s curvature, situating it naturally within the geometry of moment-restricted inference.
5.1 GMM and GEE as Quadratic Approximations
While BETEL provides the exact conditional likelihood, computational considerations often favor quadratic approximations. We show that generalized method of moments (GMM; 23) and generalized estimating equations (GEE; 43) arise naturally as optimal second-order expansions of this likelihood, with their properties governed by the same curvature parameter. This connection situates our framework within the broader theory of moment-based estimation and quasi-likelihood, linking the geometric weighting implied by $H^{*}$ to classical results on optimal GMM weighting and efficiency [46, 59, 26]. The curvature perspective thus provides a unified explanation for both efficiency under correct specification and robustness under model misspecification.
@@ 5.1 GMM and GEE as Quadratic Approximations
Thus GEE achieves robustness by replacing the exact information metric with a working approximation. Independence working correlations preserve rates whenever true correlations are moderate, while pathological choices that nearly annihilate directions in $\mathrm{span}(D_{i})$ destroy stability by driving $\lambda_{\min}(J)$ to zero.
@@ 6.1 Berk’s Theorem via Constraint Curvature
This refines the misspecified Bernstein–von Mises picture [e.g. 33]: curvature dictates the local information content and the rate at which the posterior contracts around the pseudo-true parameter. Large curvature accelerates learning yet increases brittleness under model error; small curvature slows learning but yields thicker, more robust posteriors. The same phenomenon holds predictively.
@@
Thus forecasts converge to the pseudo-true predictive law at the same curvature-controlled rate.
@@ 6.2 Safe Bayes and Geometric Tempering
Tempering therefore trades speed for robustness in a controlled, metric-level manner.
@@
Let $\widehat{\lambda}_{\min}$ be a consistent local estimate of $\lambda_{\min}(H^{*})$. Any choice $T=T(\widehat{\lambda}_{\min})$ that keeps $T/\widehat{\lambda}_{\min}$ bounded above and away from zero yields a collapse rate $O\!\big(m\sqrt{\log n/n}\big)$, up to constants uniform over models with comparable curvature. In particular, setting
@@
Implications. Berk’s theorem is not an exception to Bayesian learning but the boundary case of constrained prediction, with curvature serving as the single control parameter for both efficiency and robustness. The diagnostic is geometric (inspect $\lambda_{\min}(H^{*})$); the remedy is geometric (rescale the metric via tempering); and the consequences are transparent: fast and brittle when curvature is large, slow and robust when curvature is small, and optimally calibrated when curvature is moderated. This unifies misspecification analysis, posterior tempering, and predictive calibration within a single geometric framework. Recent results by McLatchie et al. [44] reinforce this interpretation: as the sample size grows, the temperature of a power posterior becomes asymptotically inconsequential for predictive accuracy, a limiting manifestation of predictive collapse that our curvature-adaptive theory quantifies at finite $n$.
@@ 7 Discussion and Outlook
Chernozhukov and Hong [10] reinterpret classical estimators such as GMM through quasi-posteriors, showing that standard estimators emerge as Laplace-type limits of MCMC-based Bayesian updates. Their framework uses MCMC to approximate optimal weighting and curvature. By contrast, the present analysis derives this Gaussian structure directly from first principles—exchangeability and constraint geometry—without recourse to sampling algorithms. The curvature term $\lambda_{\min}(H^{*})$ appears not as a numerical by-product but as the intrinsic geometric rate governing predictive contraction, providing a constructive explanation of the phenomena that quasi-Bayesian methods simulate.
@@ 7.1 Methodological Implications
Our analysis complements the classical Sanov route by asking the dual question: given empirical constraints, how rapidly does the predictive law return to equilibrium on the constraint manifold? This perspective makes $\lambda_{\min}(H^{*})$ the control parameter for this return, linking information geometry with predictive concentration. The Lanford window construction refines standard saddlepoint expansions by balancing lattice discretization and Gaussian localization, yielding practical non-asymptotic bounds. As the construction is coordinate invariant, the contraction rate depends only on geometric curvature, not parameterization.
@@ Acknowledgement
The initial idea for this work originated in 2014 after the authors attended a seminar on Bayesian nonparametrics and the method of moments. It was during that discussion that the possibility of an alternative route, linking predictive asymptotics to constraint geometry, first became apparent. The present framework developed gradually from that early intuition.
@@ Appendix A Proofs of Main Results
Lanford Windows and Operator–Theoretic Foundation
The Lanford window construction introduced in Section 3 admits an operator–theoretic formulation originally developed by Lanford [35, 36] for lattice gases and combinatorial systems, here adapted to constrained exchangeable sequences. The aim is to formalize the localization principle that balances Gaussian fluctuations against lattice discretization errors.
The Gaussian approximation of Theorem 3.1 is valid for $O(n^{-1/2})$ fluctuations around the information projection $P^{*}$, while the multinomial lattice introduces discretization errors of comparable order. Setting
@@
balances these two effects: empirical types farther than $\rho_{n}$ from $P^{*}$ have exponentially small weight of order $O(n^{-1})$, whereas lattice approximation errors within the window $\mathcal{W}_{n}$ remain of order $O(n^{-1/2})$.
Definition of the Lanford operator.
For functions $f:E_{n}\to\mathbb{R}$ defined on the set of feasible empirical measures, define the Lanford operator
@@
which maps test functions to their Gaussian–weighted local averages around the information projection $P^{*}$. Here $H^{*}$ denotes the information–geometric Hessian at $P^{*}$ and $\lambda_{\min}(H^{*})$ its smallest eigenvalue, quantifying local curvature.
Fixed–point characterization.
Applying $\mathcal{L}_{n}$ to the quadratic test function $f(Q)=\|Q-P^{*}\|_{2}^{2}$ gives the expected squared localization radius,
@@
This expression admits a unique fixed point satisfying
@@
and hence
$\rho_{n}=\Bigl(\frac{2\log n}{\lambda_{\min}(H^{*})}\Bigr)^{1/2}.$
The resulting window $\mathcal{W}_{n}=\{P\in E_{n}:\|P-P^{*}\|_{2}\leq\rho_{n}n^{-1/2}\}$ is therefore curvature–adaptive: it expands when the constraint manifold is flat (small $\lambda_{\min}$) and contracts under strong curvature.
Convergence of the operator iteration.
Iterating the operator $\mathcal{L}_{n}$ from any initial radius estimate $\rho_{n,0}$ yields a contractive mapping
$\rho_{n,k+1}^{2}=\mathcal{L}_{n}[\|P-P^{*}\|_{2}^{2}](P^{*}),$
which converges geometrically to the fixed point at rate $1-O(n^{-1/2})$. This self–consistent localization ensures that the Gaussian approximation remains uniformly valid within $\mathcal{W}_{n}$, providing the analytic foundation for curvature–adaptive predictive collapse.
A.1 Proof of Theorem 3.1: Tilted Finite Gaussian De Finetti
The argument localizes the predictive mixture around the information projection $P^{*}$, replacing the discrete lattice of feasible empirical types with a Gaussian integral on the tangent space $\mathsf{T}^{*}$. We decompose the constrained predictive sum into interior and exterior contributions relative to the localization window $\mathcal{W}_{n}$:
@@
where $S_{n}^{(\mathrm{in})}(A)$ aggregates empirical types $P\in\mathcal{W}_{n}$ and $S_{n}^{(\mathrm{out})}(A)$ those with $P\notin\mathcal{W}_{n}$. The exterior component collects types far from the information projection.
Tail bound.
Recall that
$\mathcal{W}_{n}=\{P\in E_{n}:\|P-P^{*}\|_{2}\leq\rho_{n}/\sqrt{n}\},\qquad\rho_{n}^{2}=\frac{2\log n}{\lambda_{\min}(H^{*})},$
where $\lambda_{\min}(H^{*})$ denotes the smallest eigenvalue of the information–geometric Hessian at $P^{*}$.
@@
For any empirical type $P\notin\mathcal{W}_{n}$,
@@
and consequently
$|S_{n}^{(\mathrm{out})}(A)|=O(n^{-1})\quad\text{uniformly for }A\subset\mathcal{X}^{m}.$
Proof of Lemma A.1.
By convexity of the Kullback–Leibler divergence,
$D_{D}(P\|Q)\geq D_{D}(P^{*}\|Q)+\tfrac{1}{2}(P-P^{*})^{\top}H^{*}(P-P^{*}).$
If $\|P-P^{*}\|_{2}>\rho_{n}/\sqrt{n}$, then $(P-P^{*})^{\top}H^{*}(P-P^{*})\geq\lambda_{\min}(H^{*})\rho_{n}^{2}/n=2\log n/n$, so the exponential term is bounded by $n^{-1}$, proving the claim. ∎
Local expansion inside $\mathcal{W}_{n}$.
For $P\in\mathcal{W}_{n}$, the log–weight admits the quadratic expansion
@@
where $P=P^{*}+n^{-1/2}v+O(n^{-1})$,
$L_{n}(v;x_{1:m})=n^{-1/2}\sum_{i=1}^{m}v^{\top}\nabla\!\log P^{*}(x_{i})$
is the linear fluctuation term, and $R_{n}(v;m)=O(\log n/n^{1/2})$ uniformly for $\|v\|\leq\rho_{n}$. The exponential factor therefore behaves as a Gaussian density on $\mathsf{T}^{*}$ with precision $nH^{*}$.
Discrete–continuous approximation.
Let $\mathcal{L}_{n}=\{n^{1/2}(P-P^{*}):P\in E_{n}\cap\mathcal{W}_{n}\}$ denote the lattice of rescaled deviations.
@@
For any function $f$ with a Gaussian envelope,
@@
Proof sketch.
The result follows from the localization technique of Lanford [35, 36], which approximates discrete lattice sums by continuous Gaussian integrals over $\mathsf{T}^{*}$, and from standard Laplace approximation techniques for lattice sums [19, 15]. Since the cell volume of the empirical lattice scales as $n^{-r/2}$ with $r=\dim(\mathsf{T}^{*})$, the discretization error is $O(n^{-1/2}\operatorname{polylog}n)$. ∎
Assembly.
Combining Lemmas A.1 and A.2 gives
$\mu_{n,m}(A)=\int_{\mathsf{T}^{*}}P_{P(v)}^{\otimes m}(A)\,\phi_{H^{*}}(v)\,dv+O\!\big(n^{-1/2}\operatorname{polylog}(n)\big),$
where $\phi_{H^{*}}(v)$ is the Gaussian density with precision matrix $H^{*}$ and $P(v)=\Pi(P^{*}+n^{-1/2}v)$ is the local chart projection onto $E$. This establishes Theorem 3.1.